OpenStudy (jjuden):

Solve: 9x^2 + 3x = 0

OpenStudy (anonymous):

factor out the x, you get x (9x + 3). From there set each equal to 0 and solve for x

OpenStudy (anonymous):

x = 0 @jjuden Do you understand why? @jjuden

OpenStudy (jjuden):

@Tom_Boy_Rebel you dont have to put @jjuden

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

there are two solutions x=0 and x=-1/3

OpenStudy (anonymous):

x = 0 or x = -1/3 do you understand why or do you need more help?

OpenStudy (jjuden):

@mathman123 i know but how do i get them

OpenStudy (anonymous):

I will show you how jjuden

OpenStudy (jjuden):

stop putting jjuden please

OpenStudy (anonymous):

use the quadratic formula to solve it \[x=\frac{ -b \pm \sqrt{b ^{2}-4ac}}{ 2a }\]

OpenStudy (jjuden):

it doesnt say to use the quadratic formula i olny use that when it tells me to

OpenStudy (agent0smith):

9x^2 + 3x = 0 Factor out an x as @alias said: x(9x + 3) = 0 now set each equal to zero: x = 0, 9x + 3 = 0 solve the one on the right, the other is already x=0

OpenStudy (anonymous):

Take the integral: sin(x) cos^2(x) dx For the integrand sin(x) cos^2(x), substitute u = cos(x) and du = -sin(x)dx: =-u^2 du The integral of u^2 is u^3/3: =-u^3/3 + constant Substitute back for u = cos(x): =-1/3

OpenStudy (agent0smith):

^why is there a random calculus solution....

OpenStudy (anonymous):

Is that related to the problem? O.O

OpenStudy (anonymous):

lol!